OpenStudy (anonymous):

how do you solve 5=\sqrt{r-3}

OpenStudy (anonymous):

you dont feather

OpenStudy (anonymous):

=> 25=r-3 => r=25+3=28

OpenStudy (anonymous):

what

OpenStudy (anonymous):

thank you anotherxx